# The Indian Pulsar Timing Array Data Release 2: II. Customised Single-Pulsar Noise Analysis and Noise Budget

### v1

@@ 3.2.1 White Noise
Each ToA measurement is associated with an inherent uncertainty due to the noise introduced by the radiometer (finite S/N of the average pulse) and the estimation techniques. Additionally, due to the presence of RFI as well as fluctuations of the single pulse profile due to processes intrinsic to the pulsar, the overall ToA uncertainty needs to be modified to get a more realistic estimate. This modification is done by accounting for white noise in the data using three parameters – (i) EFAC (Error-FACtor), which accounts for the radiometer-induced noise and under-estimated template matching errors while estimating the ToAs, (ii) EQUAD (Error added in QUADrature), which accounts for additional noise in the ToA uncertainties unmodeled by EFAC, and (iii) ECORR (Error CORRelated in radio-frequency), which quantifies the noise correlated across the observing frequencies for a particular observation but is uncorrelated across multiple observations. The latter two parameters mainly account for pulse jitter (Kulkarni+2024; Kikunaga+2024). These parameters are added for a particular set of observations based on the observing characteristics and backend-receiver configurations in this work. Both EQUAD and ECORR are added in quadrature with the ToA uncertainties such that the white noise covariance matrix can be represented as
@@
where $\sigma_{ij}$ is the uncertainty estimated by the ToA generation algorithm in the ToA for $i\text{th}$ epoch and $j\text{th}$ sub-band, $\delta_{ij}$ is the Kronecker delta, $\mathcal{U}$ is a block-diagonal matrix with $1$ for ToAs of the same observation and 0 otherwise, and the subscript $c$ refers to a particular backend-receiver configuration.
@@ 3.2.2 Achromatic Red Noise
The Achromatic Red Noise (ARN), also known as spin noise, characterises long-term, time-correlated variations in the timing residuals arising from the intrinsic irregularities in the pulsar’s spin. While ARN tends to be less pronounced in MSPs compared to young pulsars (Grover2024), it still induces slow, systematic deviations in the timing residuals of MSPs over extended timescales. We model the ARN as a stationary Gaussian Process (GP) in the Fourier domain, defined by the power-law spectral density
@@
where $A_{\text{arn}}$ is the power-law amplitude referenced to $f_{1\text{yr}}$, $\gamma_{\text{arn}}$ is the spectral index of the power-law, $f$ is the Fourier-domain frequency and $f_{1\text{yr}}$ is a characteristic reference frequency for the power-law process and corresponds to $1\,\text{yr}^{-1}$. The timing residuals due to a stochastic red noise process can be approximated as (Srivastava2023)
@@ 3.2.3 Dispersion Measure Noise
The interaction of radio pulses with the IISM introduces time-correlated, radio-frequency dependent delays in the timing residuals, giving rise to chromatic noise (Cordes1990; Kaspi1994; You2007; Petroff2013; Cordes2016; Donner2020; Goncharov+2021; Srivastava2023) due to the relative motion of the Earth, IISM and pulsar along with the dynamic nature of the IISM. In its simplest form, these effects alter the column density of electrons along the propagation path and hence the DM of the pulsar. This gives rise to additional radio-frequency dependent stochastic delays in the timing residuals which scale as
@@
These stochastic DM variations are modeled as a red noise process using a power-law Fourier basis GP which depends on both the Fourier frequency as well as radio-frequency due to the inherent chromaticity as per equation 6, with a spectral density (MPTA4.5yrNoise)
@@
where $A_{\text{dmn}}$ is the amplitude referenced to $f_{1\text{yr}}$, $\gamma_{\text{dmn}}$ is the spectral index of the power-law process and $\nu_{\text{ref}}$ is a characteristic reference radio-frequency fixed at $1400\,\text{MHz}$ in this work. This noise model is hitherto referred as the DM noise model or DMN in the rest of the paper. We have adopted a Fourier-domain GP model for DM noise rather than modeling the DM variations via DMXs in time-domain, as this approach aligns our analysis with the methodology adopted in the wider IPTA community (3P+comp).
@@ 3.2.4 Scattering Noise
The turbulent nature of the IISM leads to multi-path propagation of radio pulses through the density inhomogeneities in the plasma, causing diffractive scattering of the radio photons (LorimerKramer2004; Goncharov+2021; EPTA2021TN; EPTA+InPTA-II; Srivastava2023). This diffractive interstellar scattering manifests as time-varying broadening of the intrinsic pulse profile, thereby producing additional chromatic delays in the measured ToAs. Under the thin-screen approximation with Gaussian inhomogeneities (LorimerKramer2004), these scattering-induced delays scale as
@@
This chromatic scaling is steeper than that of the DM process, providing a way to distinguish between the two contributions. This diffractive scattering is also modeled as a chromatic red noise process using a power-law Fourier basis GP with a spectral density (MPTA4.5yrNoise)
@@ 3.2.5 Free-Chromatic Noise
The treatment discussed in Section 3.2.4 is not exhaustive since the geometry and distribution of density inhomogeneities as well as the turbulence spectrum of the IISM can vary from one line-of-sight to other. In order to take care of these effects, the chromatic index of this noise process is also set as a free parameter (hence the name ‘free’-chromatic noise), represented by $\chi_{\text{fcn}}$, such that the spectral density of the Fourier-basis GP takes the form
@@
In the present analysis, we incorporate the free-chromatic noise model instead of the scattering model to take into account the above mentioned variability in the chromatic power spectrum for different pulsars spanning different line-of-sights. This noise model is hitherto referred as FCN in the rest of the paper.
@@ 3.2.6 Solar-Wind Noise
Solar-wind is a varying, ionised stream of charged particles ejected from the Sun, which modifies the Total Electron Content (TEC) along the line-of-sight between the Earth and a pulsar (Tiburzi+2019). This variable TEC, which manifests as a variation in the DM, induces a chromatic delay in the radio signals and hence, the ToAs, in a way similar to DM variations but also depending on the pulsar’s varying solar elongation [^4] The angular separation between the pulsar and the Sun as viewed from the Earth is called the solar elongation of the pulsar. Due to the revolution of the Earth around the Sun, this elongation changes and the line-of-sight towards the pulsar appears closest to the Sun once a year. throughout the year (Tiburzi2021; Susarla2024; Liu2025). This additional chromatic delay is called the ‘solar-wind noise’ and can manifest as a chromatic red noise process, particularly dominating when the pulsar’s solar elongation is small (Susarla2024).
Following the prescription presented in Susarla2024, we model these delays in the time-domain by a deterministic model that takes care of the constant, average solar-wind electron density [^5] The average solar-wind electron density is referenced at a distance of $1$ Astronomical Unit (AU) from the Sun., represented by $n_{\text{earth}}$. Additionally, we also incorporate a first-order derivative term $\dot{n}_{\text{earth}}$ to account for the temporal variations in $n_{\text{earth}}$. The complete deterministic signal is, thus, represented as
@@
where $t_{i}$ is the observing epoch, $r_{\text{earth}}(t_{i})$ is the distance between the observatory and the Sun at $t_{i}$, $\rho(t_{i})$ is the solar elongation at $t_{i}$, $\nu_{j}$ is the observing frequency, $K_{\text{D}}\simeq 4.1488\,\text{MHz}^{2}\,\text{pc}^{-1}\,\text{cm}^{3}\,\text{s}$ (LorimerKramer2004) is the Dispersion Constant, and $t_{\text{ref}}$ is a fiducial reference epoch in the Taylor-series expansion of $n_{\text{earth}}$ and is considered as the starting epoch of observation in this work. This noise model is referred to as SWdet model hitherto in this work.
It should be noted that the delays introduced by the solar-wind effects have a similar behavior as that of a time-varying DM process, except for the modulation introduced by the varying solar elongation over the course of a year, rendering high covariance between the two processes. Preliminary analyses done including the Solar-Wind GP (SWGP) introduced in Susarla2024 along with the deterministic model showed significant spectral leakage and corruption of the DM process. While such a result may stem due to the absence of high cadence observations near solar-conjunctions (where the solar-wind effects are particularly dominant) and smaller time-baseline for many pulsars, a rigorous inspection of the SWGP model for the InPTA dataset is currently being investigated. Therefore, we are not adopting the GP treatment of solar-wind noise, rather we adopt the above outlined framework to account for the secular variations up to first order in time-domain. Apart from that, specific solar events such as the ones described in kmj+2021 and Chowdhury et al. (under review) cannot be adequately captured by this model.
@@ 3.3 Bayesian framework for Noise Modeling
Due to the complexities and inherent interplay of the plethora of noise processes outlined in Section 3.2, it is crucial to accurately characterise them, so as to adequately estimate the overall PTA noise budget. However, due to the large sample size of PTAs, it becomes computationally challenging to account for all the noise processes exhaustively for every pulsar by unnecessarily expanding the parameter space. Therefore, a careful selection and inference paradigm is required to meet the two otherwise divergent scenarios.
We perform Bayesian inference to select the optimal model for different pulsars due to its data-driven approach wherein the data informs which model it prefers the most, thereby occupying a sweet spot in PTA noise analyses where the presence of myriad signals makes it extremely complicated to assess the suitable model. The Bayes’ Theorem, in this context, can be written as
@@
The timing and noise models described in Sections 3.1 and 3.2 are incorporated in the analysis using the ENTERPRISE package (Ellis+2020; Johnson+2024). The posterior distributions of various noise processes are evaluated using the DYNESTY package (Speagle2020), which uses the nested sampling technique (Skilling2004) and the PTMCMCSampler (Ellis+2017) which employs the Parallel-Tampering Markov Chain Monte Carlo sampling technique (Metropolis+1953; Hastings1970; Earl+2005; Sanjib). We perform model selection by evaluating the Bayesian evidence from the posterior distributions as (Trotta; Sanjib; Weller; Krishak)
@@
where $\bm{r}$ corresponds to the $N$-dimensional vector of timing residuals, $\bm{\mathcal{D}}$ is the vector representing time-domain deterministic signals (such as the solar-wind noise signal in equation 11), and $\bm{C}$ is the noise covariance matrix of the stochastic processes (Susarla2024; MPTA4.5yrNoise). In general, the evaluation of the entire PTA likelihood is computationally expensive, but in the regime where the timing model perturbations can be represented by a first-order Taylor series expansion (i.e., the linear regime), the likelihood function can be analytically marginalized over the timing model parameters (Lentati2013; van_Haasteren2014). As long as we are working in the linear regime (which is true in most cases as the timing solution is optimal, given the dataset), this framework offers a computationally efficient method of likelihood implementation in PTA analyses.
In this work, we arrive at the optimal noise model and parameter estimates by incorporating a comprehensive inferencing and testing methodology which is described in the subsequent parts of this section.
@@ 3.3.1 Optimal Model Selection
This suite of models caters to the pulsar-specific S/N and jitter behavior and is different from Srivastava2023 where a sequential approach was adopted (see Section 6 for a detailed comparison). We add a white noise parameter per receiver-backend configuration for the InPTA dataset, which amounts to a maximum of 21 white noise parameters for the longest time baseline pulsars (see Section 2 and Rana+2025 for more details). The number of Fourier modes for the red noise processes is set in such a way that the highest frequency in the noise spectrum corresponds to $1\,\text{month}^{-1}$, which serves as the Nyquist cutoff for the biweekly observational cadence of InPTA (Rana+2025). The prior distributions adopted for various noise hyper-parameters are detailed in Table 1. The posterior distributions and evidences for each model are evaluated using the DYNESTY package and the Bayes Factors for different models are calculated using equation 14.
We employ the threshold of $\ln\mathcal{B}_{12}>4.6$ as per the Jeffreys scale (Jeffreys1939; Trotta; Weller) for strong statistical preference towards model 1. However, instead of accepting the strongly preferred model based on the Bayes Factor metric alone, we further investigate how well the parameter posteriors are constrained along with the post-noise residuals obtained by subtracting the noise contributions from the timing residuals. These residuals are then normalised with the scaled ToA uncertainties (using white-noise parameters) and are then subjected to the Anderson-Darling Gaussianity test (AndersonDarling1954) along with generating Q-Q plots for validation [^6] A Quantile-Quantile (Q-Q) plot is a graphical tool to compare two distributions by plotting their quantiles against each other. The standard Normal distribution is expected to follow a $45^{\circ}$ line in such a plot. These serve as visual checks for ascertaining the behavior of distributions.. In several cases, we encounter that even though the Bayes Factors strongly preferred a more complicated model, still the posterior distributions were poorly constrained and noise-subtracted residuals had a higher level of non-Gaussianity than the lesser complicated model. Such a spurious effect is termed as the prior volume effect in literature (see Gomez+2022 and the references therein) and mainly arises due to the sensitivity of the integral in equation 13 to the volume of the prior distribution in the parameter space. In case of PTA analyses, this effect is more susceptible to occur due to the use of uninformed priors for noise hyper-parameters.

### v2

@@ 3.2.1 White Noise
Each ToA measurement is associated with an inherent uncertainty due to the noise introduced by the radiometer (finite S/N of the average pulse) and the estimation techniques. Additionally, due to the presence of RFI as well as fluctuations of the single pulse profile due to processes intrinsic to the pulsar, the overall ToA uncertainty needs to be modified to get a more realistic estimate. This modification is done by accounting for white noise in the data using three parameters – (i) EFAC (Error-FACtor), which accounts for the radiometer-induced noise and under-estimated template matching errors while estimating the ToAs, (ii) EQUAD (Error added in QUADrature), which accounts for additional noise in the ToA uncertainties unmodeled by EFAC, and (iii) ECORR (Error CORRelated in radio-frequency), which quantifies the noise correlated across the observing frequencies for a particular observation but is uncorrelated across multiple observations. The latter two parameters mainly account for pulse jitter (Kulkarni et al., 2024; Kikunaga et al., 2024). These parameters are added for a particular set of observations based on the observing characteristics and backend-receiver configurations in this work. Both EQUAD and ECORR are added in quadrature with the ToA uncertainties such that the white noise covariance matrix can be represented as
@@
where $i,j$ denote the ToA indices, $\sigma$ is the uncertainty estimated by the ToA generation algorithm, $\delta_{ij}$ is the Kronecker delta, $\mathcal{U}$ is a block-diagonal matrix with $1$ for ToAs of the same observation and 0 otherwise, and the subscript $c$ refers to a particular backend-receiver configuration.
@@ 3.2.2 Achromatic Red Noise
The Achromatic Red Noise (ARN), also known as spin noise, characterizes long-term, time-correlated variations in the timing residuals arising from the intrinsic irregularities in the pulsar’s spin. While ARN tends to be less pronounced in MSPs compared to young pulsars (Grover et al., 2024), it still induces slow, systematic deviations in the timing residuals of MSPs over extended timescales. We model the ARN as a stationary Gaussian Process (GP) in the Fourier domain, defined by the power-law spectral density
@@
where $A_{\text{arn}}$ is the power-law amplitude referenced to $f_{1\text{yr}}$, $\gamma_{\text{arn}}$ is the spectral index of the power-law, $f$ is the Fourier-domain frequency and $f_{1\text{yr}}$ is a characteristic reference frequency for the power-law process and corresponds to $1\,\text{yr}^{-1}$. The timing residuals due to a stochastic red noise process can be approximated as (Srivastava et al., 2023)
@@ 3.2.3 Dispersion Measure Noise
The interaction of radio pulses with the IISM introduces time-correlated, radio-frequency dependent delays in the timing residuals, giving rise to chromatic noise (Cordes et al., 1990; Kaspi et al., 1994; You et al., 2007; Petroff et al., 2013; Cordes et al., 2016; Donner et al., 2020; Goncharov et al., 2021a; Srivastava et al., 2023) due to the relative motion of the Earth, IISM and pulsar along with the dynamic nature of the IISM. In its simplest form, these effects alter the column density of electrons along the propagation path and hence the DM of the pulsar. This gives rise to additional radio-frequency dependent stochastic delays in the timing residuals which scale as
@@
These stochastic DM variations are modeled as a red noise process using a power-law Fourier basis GP which depends on both the Fourier frequency as well as radio-frequency due to the inherent chromaticity as per equation 6, with a spectral density (Miles et al., 2025a)
@@
where $A_{\text{dmn}}$ is the amplitude referenced to $f_{1\text{yr}}$, $\gamma_{\text{dmn}}$ is the spectral index of the power-law process and $\nu_{\text{ref}}$ is a characteristic reference radio-frequency fixed at $1400\,\text{MHz}$ in this work. This noise model is henceforth referred as the DM noise model or DMN in the rest of the paper. We have adopted a Fourier-domain GP model for DM noise rather than modeling the DM variations via DMXs in time-domain, as this approach aligns our analysis with the methodology adopted in the wider IPTA community (Agazie et al., 2024).
@@ 3.2.4 Scattering Noise
The turbulent nature of the IISM leads to multi-path propagation of radio pulses through the density inhomogeneities in the plasma, causing diffractive scattering of the radio photons (Lorimer and Kramer, 2004; Levin et al., 2016; Chalumeau et al., 2022; EPTA Collaboration et al., 2023; Turner et al., 2021; Liu et al., 2022; Srivastava et al., 2023). This diffractive interstellar scattering manifests as time-varying broadening of the intrinsic pulse profile, thereby producing additional chromatic delays in the measured ToAs. Under the thin-screen approximation with Gaussian inhomogeneities (Lorimer and Kramer, 2004), these scattering-induced delays scale as
@@
This chromatic scaling is steeper than that of the DM process, providing, in principle, a way to distinguish between the two contributions. However, because of significant measurement uncertainties and limited (fractional) bandwidth, the $\nu^{-2}$ and $\nu^{-4}$ scalings appear similar, in practice, making it difficult to distinguish between the two processes. This was recently demonstrated on EPTA data (Iraci et al., 2025). This diffractive scattering is also modeled as a chromatic red noise process using a power-law Fourier basis GP with a spectral density (Miles et al., 2025a)
@@ 3.2.5 Free-Chromatic Noise
The treatment discussed in Section 3.2.4 is not exhaustive since the geometry and distribution of density inhomogeneities as well as the turbulence spectrum of the IISM can vary from one line-of-sight to another. In order to take care of these effects, the chromatic index of this noise process is also set as a free parameter (hence the name ‘free’-chromatic noise), represented by $\chi_{\text{fcn}}$, such that the spectral density of the Fourier-basis GP takes the form
@@
In the present analysis, we incorporate the free-chromatic noise model instead of the scattering model to take into account the above mentioned variability in the chromatic power spectrum for different pulsars spanning different line-of-sights. This noise model is henceforth referred to as FCN in the rest of the paper.
@@ 3.2.6 Solar-Wind Noise
Solar-wind is a varying, ionized stream of charged particles ejected from the Sun, which modifies the Total Electron Content (TEC) along the line-of-sight between the Earth and a pulsar (Tiburzi et al., 2019). This variable TEC, which manifests as a variation in the DM, induces a chromatic delay in the radio signals and hence, the ToAs, in a way similar to DM variations but also depending on the pulsar’s varying solar elongation  [^5] The angular separation between the pulsar and the Sun as viewed from the Earth is called the solar elongation of the pulsar. Due to the revolution of the Earth around the Sun, this elongation changes and the line-of-sight towards the pulsar appears closest to the Sun once a year. throughout the year (Tiburzi et al., 2021; Susarla et al., 2024; Liu et al., 2025). This additional chromatic delay is called the ‘solar-wind noise’ and can manifest as a chromatic noise process, particularly dominating when the pulsar’s solar elongation is small (Susarla et al., 2024).
Following the prescription presented in Susarla et al. (2024), we model these delays in the time-domain by a deterministic model that takes care of the constant, average solar-wind electron density [^6] The average solar-wind electron density is referenced at a distance of $1$ Astronomical Unit (AU) from the Sun., represented by $n_{\text{earth}}$. Additionally, we also incorporate a first-order derivative term $\dot{n}_{\text{earth}}$ to account for the temporal variations in $n_{\text{earth}}$ [^7] The long-term variations in the solar-wind electron density mimic the periodicity of the 11-year solar-cycle. The time-span of our dataset which is most sensitive to such variations is only $\sim$$3.5$ years. Thus, a deterministic process up to first order derivative term is sufficient to model such variations.. The complete deterministic signal is thus, represented as
@@
where $t_{i}$ is the observing epoch, $r_{\text{earth}}(t_{i})$ is the distance between the observatory and the Sun at $t_{i}$, $\rho(t_{i})$ is the solar elongation at $t_{i}$, $\nu_{j}$ is the observing frequency, $K_{\text{D}}\simeq 4.1488\,\text{MHz}^{2}\,\text{pc}^{-1}\,\text{cm}^{3}\,\text{s}$ (Lorimer and Kramer, 2004) is the Dispersion Constant, and $t_{\text{ref}}$ is a fiducial reference epoch in the Taylor-series expansion of $n_{\text{earth}}$ and is considered as the starting epoch of observation in this work. This noise model is henceforth referred to as SWdet model in this work.
It should be noted that the delays introduced by the solar-wind effects have a similar behavior as that of a time-varying DM process, except for the modulation introduced by the varying solar elongation over the course of a year, rendering high covariance between the two processes. Preliminary analyses done including the Solar-Wind GP (SWGP) introduced in Susarla et al. (2024) along with the deterministic model showed significant spectral leakage and corruption of the DM process. While such a result may stem from the absence of high cadence observations near solar-conjunctions (where the solar-wind effects are particularly dominant) and smaller time-baseline for many pulsars, a rigorous inspection of the SWGP model for the InPTA dataset is currently being investigated. Therefore, we are not adopting the GP treatment of solar-wind noise, rather we adopt the above outlined framework to account for the secular variations up to first order in time-domain. Apart from that, specific solar events such as the ones described in Krishnakumar et al. (2021) and  Chowdhury et al. (2026) cannot be adequately captured by this model.
@@ 3.3 Bayesian framework for Noise Modeling
Due to the complexities and inherent interplay of the plethora of noise processes outlined in Section 3.2, it is crucial to accurately characterize them, so as to adequately estimate the overall PTA noise budget. However, due to the large sample size of PTAs, it becomes computationally challenging to account for all the noise processes exhaustively for every pulsar by unnecessarily expanding the parameter space. Therefore, a careful selection and inference paradigm is required to meet the two otherwise divergent scenarios.
We perform Bayesian inference to select the optimal model for different pulsars due to its data-driven approach wherein the data informs which model it prefers the most, thereby occupying a sweet spot in PTA noise analyses, where the presence of myriad signals makes it extremely complicated to assess the suitable model. The Bayes’ Theorem, in this context, can be written as
@@
The timing and noise models described in Sections 3.1 and 3.2 are incorporated in the analysis using the ENTERPRISE package (Ellis et al., 2020; Johnson et al., 2024). The posterior distributions of various noise processes are evaluated using the DYNESTY package (Speagle, 2020), which uses the nested sampling technique (Skilling, 2004) and the PTMCMCSampler (Ellis and van Haasteren, 2017) which employs the Parallel-Tampering Markov Chain Monte Carlo sampling technique (Metropolis et al., 1953; Hastings, 1970; Earl and Deem, 2005; Sharma, 2017). We perform model selection by evaluating the Bayesian evidence from the posterior distributions as (Trotta, 2017; Sharma, 2017; Kerscher and Weller, 2019; Krishak and Desai, 2020)
@@
where $\bm{r}$ corresponds to the $N$-dimensional vector of timing residuals, $\bm{\mathcal{D}}$ is the vector representing time-domain deterministic signals (such as the solar-wind noise signal in equation 11), and $\bm{C}$ is the noise covariance matrix of the noise processes (Susarla et al., 2024; Miles et al., 2025a). In general, the evaluation of the entire PTA likelihood is computationally expensive, but in the regime where the timing model perturbations can be represented by a first-order Taylor series expansion (i.e., the linear regime), the likelihood function can be analytically marginalized over the timing model parameters (Lentati et al., 2013; van Haasteren and Vallisneri, 2014). As long as we are working in the linear regime (which is true in most cases as the timing solution is optimal, given the dataset), this framework offers a computationally efficient method of likelihood implementation in PTA analyses.
In this work, we arrive at the optimal noise model and parameter estimates by incorporating a comprehensive inferencing and testing methodology, which is described in the subsequent parts of this section.
@@ 3.3.1 Optimal Model Selection
This suite of models caters to the pulsar-specific S/N and jitter behavior and is different from Srivastava et al. (2023) where a sequential approach was adopted (see Section 6 for a detailed comparison). We add a white noise parameter per receiver-backend configuration for the InPTA dataset, which amounts to a maximum of 21 white noise parameters for the longest time baseline pulsars (see Section 2 and Rana et al. (2025a) for more details). The number of Fourier modes for the red noise processes is set in such a way that the highest frequency in the noise spectrum corresponds to $1\,\text{month}^{-1}$, which serves as the Nyquist cutoff for the biweekly observational cadence of InPTA (Rana et al., 2025a). The prior distributions adopted for various noise hyper-parameters are detailed in Table 1. The posterior distributions and evidences for each model are evaluated using the DYNESTY package and the Bayes Factors for different models are calculated using equation 14.
We employ the threshold of $\ln\mathcal{B}_{12}>4.6$ as per the Jeffreys scale (Jeffreys, 1939; Trotta, 2017; Kerscher and Weller, 2019) for strong statistical preference towards model 1. However, instead of accepting the strongly preferred model based on the Bayes Factor metric alone, we further investigate how well the parameter posteriors are constrained along with the post-noise residuals obtained by subtracting the noise contributions from the timing residuals. These residuals are then normalized with the rescaled ToA uncertainties (using white-noise parameters) and are then subjected to the Anderson-Darling Gaussianity test (Anderson and Darling, 1954) along with generating Q-Q plots for validation [^8] A Quantile-Quantile (Q-Q) plot is a graphical tool to compare two distributions by plotting their quantiles against each other. The standard Normal distribution is expected to follow a $45^{\circ}$ line in such a plot. These serve as visual checks for ascertaining the behavior of distributions.. In several cases, we find that even though the Bayes Factors strongly preferred a more complicated model, still the posterior distributions were poorly constrained and noise-subtracted residuals had a higher level of non-Gaussianity than the less complicated model. Such a spurious effect is termed as the prior volume effect in literature (see Gómez-Valent (2022) and references therein), and mainly arises due to the sensitivity of the integral in equation 13 to the volume of the prior distribution in the parameter space. In case of PTA analyses, this effect is more susceptible to occur due to the use of uninformed priors for noise hyper-parameters.